OpenStudy (anonymous):

limit as x approaches 3 of (1+(1/3)x) =2

OpenStudy (anonymous):

Do you have a question about the limit?

OpenStudy (anonymous):

delta and epsilon

OpenStudy (anonymous):

i basically need to show proof

OpenStudy (anonymous):

What do you have so far?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3}1+\frac{ 1 }{ 3 }x=2\]Is this the limit you're talking about?

OpenStudy (anonymous):

3-3e<x<3e+3

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

You went in the wrong direction You started with abs(1+x/3-2)<e = abs(x/3-1)<e Mutliply through by 3 you have abs(x-3)<3e. See how this is related to your delta

OpenStudy (anonymous):

We prove the limit through substitution. Substitute 3 in for x in the limit.\[\lim_{x \rightarrow 3}1+\frac{ 1 }{ 3 }x=1+\frac{ 1 }{ 3 }(3)=1+1=2\]Therefore, as x approaches 3, the limit approaches 2.

OpenStudy (anonymous):

@ehhhlalaxd

OpenStudy (anonymous):

how would you prove it showin epsilon and delta

OpenStudy (anonymous):

ahhh i see @Xavier

OpenStudy (anonymous):

Oh you are talking about delta epsilon proofs of limits, no wonder you were using absolute value and stuff. I would show but I'm too tired at the moment and going to sleep. You can search up delta-epsilon proofs of limits on google, there is some easy to follow guides. Ciao.

OpenStudy (anonymous):

thank you too genius!